OpenStudy (anonymous):

How to prove that \sqrt(2)*a+\sqrt(3)*b for any non-zero integers a abd b is irational?

OpenStudy (zarkon):

prove by contradiction

OpenStudy (zarkon):

assume \[\sqrt{2}a+\sqrt{3}b=\frac{m}{n}\] where \(m,n\in\mathbb{N}\) and derive a contadiction.